# Can Alice do science and have friends, in a relational quantum world? Solipsism and Relational Quantum Mechanics

### v1

@@ Abstract
Some authors have stated that relational understandings of quantum phenomena [Kochen 1985, Bene-Dieks 2002, Berkovitz-Hemmo 2006, Conroy 2012, Mermin-Fuchs-Schack 2014, Healey 2012, Brukner 2015, Auffèves-Grangier, 2016 Zwirn, 2016] and in particular Relational Quantum Mechanics (RQM) [Rovelli 1996], lead to a form of solipsism [Pienar 2021], and worried that this solipsism could undermine the possibility of doing science.
@@
In this note, I inquire whether or not this line of reasoning represents a problem for RQM. The specific question I address is whether RQM undermines an agent’s possibility of doing science collectively with other agents, learning, and relying upon, science. I argue here that this is definitely not the case, hence these worries are unfounded. In a relational quantum world, agents like us can do science, do it collectively, and be friends.
@@
Riedel argues that this perspectivalism may lead to a sort of infinity regress. But it is so only if we anxiously search for an ultimate perspective. To understand the world, ourselves, our science, and the others, we do not need such ultimate perspective. We can simply be happy of always being, talking and describing the state of the world from within a perspective. There is no arrogance of having a privileged view here, precisely because common sense always tells us that our perspective (on contingent facts, not about laws of nature) is one among many similar others. This is true regarding motion versus stasis, it can also be true regarding values of physical variable.
@@
Thanks to Huw Price and Claudio Calosi for useful exchanges and in particular to Andrea Di Biagio and Tim Riedel for very insightful reactions to early drafts of this note.

### v2

@@ Abstract
Some authors have stated that relational understandings of quantum phenomena [Kochen 1985, Bene-Dieks 2002, Berkovitz-Hemmo 2006, Conroy 2012, Mermin-Fuchs-Schack 2014, Healey 2012, Brukner 2015, Auffèves-Grangier, 2016 Zwirn, 2016, Cavalcanti 2021] and in particular Relational Quantum Mechanics (RQM) [Rovelli 1996], lead to a form of solipsism [Pienar 2021], and worried that this solipsism could undermine the possibility of doing science.
@@
In this note, I inquire whether or not this line of reasoning represents a problem for RQM. The specific question I address is whether RQM undermines an agent’s possibility of doing science collectively with other agents, learning, and relying upon, science. I argue here that this is definitely not the case, hence these worries are unfounded. In a relational quantum world, agents like us can do science, do it collectively, and be friends. (See also [Cavalcanti 2021].)
@@
A hard-core realist can still postulate that all relative events are absolute facts. It suffices to carefully distinguish absolute facts from the relative facts that can be used as inputs for the transition amplitudes giving the probabilities or events relative to a single system. The price to pay for holding this hard-core realism is that absolute reality is going to strongly violate intuitive aspects of our common-sense spacetime reality [Bong et al 2020]. The advantage of this position, which is dear to many, is not clear to me: what is the point of postulating something that is in principle inaccessible and is there only to reproduce something that played a role in an earlier stage of lesser knowledge?
Riedel argues that such perspectivalism may lead to a sort of infinity regress. But it is so only if we anxiously search for an ultimate perspective. To understand the world, ourselves, our science, and the others, we do not need such ultimate perspective. We can simply be happy of always being, talking and describing the state of the world from within a perspective. There is no arrogance of having a privileged view here, precisely because common sense always tells us that our perspective (on contingent facts, not about laws of nature) is one among many similar others. This is true regarding motion versus stasis, it can also be true regarding values of physical variable.
@@
Thanks to Eric Cavalcanti, Emily Adlam for extended discussions on the issues discussed in this note. To Huw Price and Claudio Calosi for useful exchanges and in particular to Andrea Di Biagio and Tim Riedel for very insightful reactions to early drafts of this note.